# Eigenweights for arithmetic Hirzebruch Proportionality

### v1

@@ Abstract.
All the mathematical content of this paper was fully generated by a math-research agent powered by Gemini Deep Think, internally codenamed Aletheia at Google DeepMind (see §2). Aside from building Aletheia, the author’s contribution was only to rewrite the mathematical content into paper form.
@@ 1.3.7. Type D
(2) If $n=2m$ is even, then for $k\neq m$ the eigenweight attached to $p_{k}^{(2)}$ is the same expression (LABEL:eq:D-eps-k). The elements $p_{m}^{(2)}$ and $\operatorname{Pf}$ are not eigenvectors for $\overline{\nabla}^{\eta}_{\mu}$, but we calculate the action of $\overline{\nabla}^{\eta}_{\mu}$ on their 2-dimensional span, in (5.2.14).
@@ 2. Declaration of AI Usage
The mathematical content of this paper was entirely generated [^4] For transparency, the original prompts and model responses can be found at this link., end-to-end, by an internal reasoning agent built upon Gemini Deep Think, codenamed Aletheia at Google DeepMind [aletheia]. The paper itself was rewritten by the human author, starting from Aletheia’s output. However, by comparing the paper to the model outputs linked above, the reader can verify that Aletheia’s original output is already complete and correct in mathematical content. Thus the author’s only contribution to this paper (aside from developing Aletheia) was to fiddle with the exposition, and to write an Introduction.
Human-AI Interaction Card  Query eigenweights for Type A. Human Aletheia Response (fully correct). Proceed to query eigenweights for Type C. Human Aletheia Response (fully correct). Proceed to query eigenweights for Type D. Human Aletheia Response (fully correct). Raw prompts and outputs:  https://github.com/google-deepmind/superhuman/tree/main/aletheia.
@@ 3.1.3. Casimir
For $k=1,\ldots,n$, we have555even for $k=1$, since $p_{0}(A)=m$
@@
In these terms, we can rewrite $p_{1}(A)^{N}p_{k-1}(A)=\Lambda_{k}p_{\nu_{k}}(A)$, where $\nu_{k}$ is the partition666for $k=1$, we interpret this as $\nu_{1}=(1^{N})$. $(k-1,1^{N})$ and

### v2

@@ Abstract.
The mathematical content of this paper was fully generated by a math-research agent powered by Gemini Deep Think, internally codenamed Aletheia at Google DeepMind (see §2). Aside from building Aletheia, the author’s contribution was only to rewrite the mathematical content into paper form.
@@ 1.3.7. Type D
(2) If $n=2m$ is even, then for $k\neq m$ the eigenweight attached to $p_{k}^{(2)}$ is the same expression (1.3.6). The elements $p_{m}^{(2)}$ and $\operatorname{Pf}$ are not eigenvectors for $\overline{\nabla}^{\eta}_{\mu}$, but we calculate the action of $\overline{\nabla}^{\eta}_{\mu}$ on their 2-dimensional span, in (5.2.14).
@@ 2. Declaration of AI Usage
The core mathematical content of this paper was entirely generated by an internal reasoning agent built upon Gemini Deep Think, codenamed Aletheia at Google DeepMind [FTB+]. The paper itself was (re)written by the human author, starting from Aletheia’s output. The “Human–AI Interaction Card” (see [FTB+, §5] for explanation of the concept) below this paragraph summarizes the interaction, and links to the raw model outputs. By comparing the paper to the original outputs, the reader can verify that Aletheia’s original output is already complete and correct in mathematical content. Thus the author’s only contribution to this paper (aside from developing Aletheia) was to fiddle with the exposition, and to write an Introduction.
Human-AI Interaction Card  Query eigenweights for Type A. Human Aletheia Response (fully correct). Proceed to query eigenweights for Type C. Human Aletheia Response (fully correct). Proceed to query eigenweights for Type D. Human Aletheia Response (fully correct). Raw prompts and outputs:  https://github.com/google-deepmind/superhuman/tree/main/aletheia/F26
@@ 3.1.3. Casimir
For $k=1,\ldots,n$, we have444even for $k=1$, since $p_{0}(A)=m$
@@
In these terms, we can rewrite $p_{1}(A)^{N}p_{k-1}(A)=\Lambda_{k}p_{\nu_{k}}(A)$, where $\nu_{k}$ is the partition555for $k=1$, we interpret this as $\nu_{1}=(1^{N})$. $(k-1,1^{N})$ and